UT Sep 06 2022 - Night #5 of Restart




Night Info

  • Sunset 18:39
  • Nautical Twilight Ends 19:33
  • Astro Twilight Ends 20:03
  • Astro Twilight Begins 04:33
  • Nautical Twilight Begins 05:03
  • Sunrise 05:58
  • Moonset 02:40
  • Moon illumination ~77%

Staff

  • OSA: D.G. Huerta
  • ISp: E.Solheid
  • Mtn Manager: B. Waggoner
  • ISA/SO: B.Rothberg, A.Becker, R.T.Gatto
  • Technical Expert: J.Hill

Night Plan

Starting Config LUCI-Bino

Telescope Wakeup

  • DONE Field Aberration (2 hrs) - Critical (1hour) (SX more important than DX due to SX removal)
    • Can be done in parallel (data collection, analysis in serial)
    • need both AGW1 and AGW2 on the telescope and checked to execute
    • must complete SH spot centering and hotspot check before executing
    • 1.2" or better seeing (1" or better ideal)
    • Target Fields:
      • DONE BS9177 18:39:33.7 +49:05:38.0
      • StoneP1 22:37:45.0 -01:13:40.45
      • BS9188 23:30:33.4 +38:18:57.0
      • BS9107 03:32:03.0 +37:20:40.0

LUCI Wakeup

  • Pending LUCI issues:
    • Wavelength calibration
    • LUCI-1 no MOS masks
    • Some flexure (spectroscopy) stll pending
  • LUCI Scripts Location: Service_Observing/Restart_2022/LUCI
    • LUCI_2022_Restart_V1.77.xml - all scripts (except LS Rot ZP)
    • RotZP _v2.xml - for LS RotZP scripts
  • DONE Rotator center
  • rotator zeropoint (requires MOS mask) - (0.5 hr)
    • 1.2" or better seeing needed
    • LUCI-1 is empty - use the LS method
    • DONE LUCI-2 - tentatively OK - the last two masks are consistent with < 0.1, but the last mask was done in poor seeing
    • Masks:
      • ID 903033 - Stone O
      • ID 921934 - Stone A
      • iD 931027 - Stone D
      • Do the LS method Bino to compare L2 with mask and LS method
    • Do a MOS acquisition (imaging of mask and thru-mask image confimration) - verify the theta < 0.1 deg for each mask
    • To update results (if needed)
      • ssh -X into TO1 as telescope; then ssh - X as tcs@tcs1; The file location is in /home/telescope/TCS/Configuration/PCS/PCSInstrument.conf
    • Longslit Method (backup in case mask exchange(s) aren't completed in time
      • requires target field with two sources in slit, use alignment holes in longslit. Need LMS file with source in longslit (preferably two) - astrometric field ideal
      • use pseudo-LMS files below for acquisition - works in similar manner to MOS acquisition - will need to manually load the LMS files in RTD
  • Collect transform data - (2hrs) - must do RotZP First
    • 1.2" or better seeing needed
    • Target Fields - select one best visible field
      • 21 37 31 +00 31 55 Stone_O_m12.list
      • 22 37 45 +01 31 33 Stone_P1.list
      • 00 54 02 -00 32 34 Stone_A3.list
    • Note the LUCI frame numbers associated with each offset in IDL script
  • Imaging check (1hr)
    • ANY seeing OK
    • dither in-field
      • NGC 7027 - cycles through all filters (4pt dither) - visible until 4am
    • dither out-of-field (run one)
      • M52 - JHK, dither 5' out of field - visible all night
      • M40 - Broad filters, few narrow filters - visible all night
  • Non-sidereal imaging check (1hr)
    • dither in-field
      • Comet 22P/Kopoff - JHK - 0.5hr
        • RA 01 40 Dec +02 44 - visible 23:30 - twilight - two scripts both should work for before UT 08hr and up to UT 12hr
    • DONE dither out-of-field
      • Saturn - H2, FeII, HeI - 0.5hr
        • RA 21hr -16 Dec - visible UT 04hr-08hr - script works for entire visibility window
  • DONE LUCI Photometric Standards
    • Any conditions - must be clear
    • File Location: /home/lbto/Service_Observing/2022A_D_E/PoorWeather/LUCI_Phot_Standards/LUCI_Phot_ZP_OT_V5.xml
    • Pick any standard thats visbile, please try to get 2 diferent airmasses

LBC Checkout Continued

  • Superfocus sequence -PA=0,180, 90, 270 - high (1 hr)
    • Requires seeing 1.2" or better, some clouds (thicker clouds require brighter stars)
  • Standard photometric fields - medium - (2hr)
    • must be clear, any seeing
    • minimum two airmasses needed
  • Spectro-photometric standards - low - (1hr)
    • must be clear, any seeing


Night Log

Closed Dome

  • UT 01:11 - bringing up LUCI GUIs
  • UT 01:12 - running init all, OK , moving to N375 on LUCI-2
  • UT 01:13 - taking test darks
    • L1 0325-0326 - 2.51sec dark
    • L2 0300-0301 - 2.51sec dark (amplifier 2 still shows lower gains)
  • UT 01:22 - O.Kuhn running LBC tests on filter motions
  • UT 01:38 - LBCs released for night use
  • UT 01:49 - dome open
  • Had to set up the data subdirectory in /home/lbto to run WFS stuff, since apparently its NOT network mounted (unlike the old obs machines and ignoring what SciOps has been requesting/pleading for 2 years)
  • hopefully the wfsc idl routines won't bomb on obs2 (I run them on OBS1 last night

Field Aberration Data Collection

  • UT 01:51 - adding in values of tip and tilt from last night pupil wobble - and got an error on DX M3
    • one of the legs is having an issue
  • UT 02:04 - still too bright for pointing star
  • UT 02:14 - at collimation star (that was fast, no need to rush tonight)
  • UT 02:20 - waiting a bit for skies to darken and seeing to settle down
  • UT 02:20 - updating elevation limits
  • UT 02:23 - send an active preset to BS9177 on-axis
  • UT 02:26 - sending a guide presetn on-axis to BS9177
  • UT 02:38 - staring field_collect for SX and DX - guider shows ~1" seeing
  • UT 02:41 - restarting with limit=350, not sure how many we got before constant unable to collimate below 250 rms
  • UT 02:55 - seeing seems more steady now
  • UT 04:01 - field_collect limit switch doesn't seem to be valid for >300, setitng limit=350 still goes a popup if its >300 but < 350
  • UT 04:02 - SX is finished, DX is still going
  • UT 04:34 - DX finished
  • Saved on OBS2

LUCI Rotator Center

  • UT 04:37 - M52 - PA0 Active Preset
    • L1 0327-0328, L2 0301-0302 (2nd image is sky)
  • UT 04:45 - M52 PA 180 Track Preset
    • L1 0329, L2 0303
    • LUCI-2 out of focus?
    • Initializing detector focus on LUCI-2
    • taking test image 2.51sec
    • moving to N1.8 camera and then back - detector focus moved -244
    • moving camera back to N375, detector focus is at -312
    • taking test exposure L2 0306 - still out of focus
    • move to N1.8 and take test image (detector focus -244), N1.8 also looks out of focus
    • Sending the active preset again, to get us collimated on LUCI-2 AGW2 just in case this is casued by that (although the guidestar looked GREAT in the guidercam)
    • LUCI2 MCU says detector focus -312, collimation in AGW2 looks great, taking exposure L308
  • UT 05:00 - sending the Track PA 180 again - see if its ok - could it be that the offset to sky messed something up on the AGW side?
    • L1 0329, L2 0309
    • L1 looks good, L2 becomes elongated in each NDIT
  • UT 05:02 - sending Track PA 90
    • L1 0330, L2 0310
    • L1 looks good, L2 out of focus and elongated
  • UT 05:07 - making a Active version of PA 270 - we'll see if rotating 90deg and having active collimation - saving as a CUBE so its visisble what happens
    • L1 0331, L2 0311 - looks good
  • UT 05:13 - Active preset version of PA 90 - to see if it keeps it in focus - bad GS
  • UT 05:16 - Active preset version of PA 90
    • now says waiting for collimation on LUCi-2, took a while before wait for collimation could be aborted
    • resending again
    • not sure what to make of it
  • UT 05:23 - doing a TRACK preset for PA 270, then will do TRACK for 90 - saved as a cube
    • L1 0334 , L2 0312 - PA 270 - saved as integrated
    • L1 0335, L2 0313 -PA 90 - saved as CUBE
    • OK,didn't do it again...
  • Will postpone RotZP until next night

Rotator ZP

  • StoneO bino script 1" LS
  • Couldn't find pseudo-LMS files because they aren't on OBS2, only OBS1. So now when we upload scripts we have to manually copy to two different OBS machines every time (three if we use ROBS). This is NOT acceptable for science operations.
  • Unable to do LS method - get error message
    • Failed to init MOS GUI. IllegalArgumentException: Unsusual number of slits (10) in LMS file
    • L1 0336-0338
    • L2 0315-0317
    • Contacted G.Eychaner and J.Power
  • UT 05:47 - authorizing LUCI-2 only - will do RotZP in Stone O with MOS mask
    • L2 0317-0322
    • eliminating #36, Err 0.068,0.138
    • shifts x=1.481", y=-3.490", theta =-0.134
    • removing #33 only changed delta theta=-0.12
  • UT 06:17 - Stone A MOS mask LUCI-2
    • Guiders shows 0.7" seeing
    • all stars good Err 0.056, 0.057
    • Xshift = 1.737", Yshift = -3.462", theta=-0.079"
    • L2 0323-0326
  • Theta is on the border for LUCI-2, would like to get another datapoint (Stone D later in the night), LUCI-1 is on hold

LUCI Imaging Checkout

  • UT 06:34 - Saturn NS imaging - use 699 in NSIGUI
  • UT 06:41 - had to update, JPL horizon that I used didn't match NSIGUI output
  • UT 06:42 - sending preset
  • UT 06:44 - no GS found on AGW1 but AGW2 is happy
  • Weird IIF error - scripts were sent as Active but NonSideral
  • Tue Sep 6 06:46:01.945 2022 left Preset command rejected for LUCIFER because mixed tracking mode Presets unsupported
    Tue Sep 6 06:46:01.945 2022 right Preset command rejected for LUCIFER because mixed tracking mode Presets unsupported
  • Tue Sep 6 06:48:40.423 2022 right Preset command rejected for LUCIFER because duplicate right synchronous PresetTue Sep 6 06:48:41.132 2022 left BinocularControl command started for LUCIFER with flag SYNCPRESETTue Sep 6 06:48:41.148 2022 left BinocularControl alert for LUCIFER: BinocularControl SYNCPRESET command ignored: Duplicate requestTue Sep 6 06:48:41.166 2022 left BinocularControl command complete with warnings for LUCIFER with flag SYNCPRESETTue Sep 6 06:48:41.769 2022 left Preset command rejected for LUCIFER because duplicate left synchronous Preset
  • IIF seemed to retain some NS parameters
  • UT 06:50 - doing a pointing check now
  • UT 06:55 - GS is a double star and its centroidinig between the two
  • GS_Double.jpg
    • Selecting a new GS (it did NOT look like a double in the OT)
    • L1 0339-0339
    • L2 0327-0328
    • forgot to pause while making new script but still looks cool
  • New guisestar looks fainter but >>1mag
  • UT 07:04 - Trying 3rd GS
  • Seeing was 1.1" ast start- but slowly blowing up
  • UT 07:29 - highly variable seeing now
  • HeI image (L1 0356, L2 0344 is totaly saturated)
  • Trying OH 1060 - 2.51 sec (L1 0357)
  • Trying OH1060+z - 2.51sec (L1 0358) / and put blinds in L2
  • Seeing is 2.2" on guiders so don't feel bad dear reader...
    • continuing script with crossed filters
    • Weird L2 is significantly brighter! But L2 dichroic cuts off all flux <0.95 microns
  • UT 07:52 - trying H2+ND2 L1 0366 - looks good - no
    • L1 0367-0373
    • L2 0352-0358
  • UT 07:57 - dear reader, the guiders show 1.4"-2.2" seeing
  • UT 08:08 - now says 3.35" on the guiders...
  • But in all seriousness - tested NS guiding in-field and large offsets for > 1 hour - no issues
    • offset worked AGW1 but star so poor in seeing got anerror star verification failed
  • L1 0340-0377
  • L2 0329-0362

LUCI Photometric Standards

  • Given how poor the seeing is, going to do Phot Standards (since its clear)
  • UT 08:16 - sending preset to GSPC 893-D Airmass 1.1-1.2
  • UT 08:23 - guidestar vanished on -20,20 offset but only on LUCI-1, OK on LUCI-2
    • telsvc 0,0 absolute offset recvoered guidestar, and then 20,-20 offset worked
    • now fails +20,+20
    • L1 0378-0381
    • L2 0363-0366
    • IT 8495 - AGW1 doesn't find star but AGW2 does - see IT for position of star in patrol field
    • picked a new PA and guidestar
  • UT 08:37 - new script starting
  • UT 08:52 - Guiders show 1.8"-2.2" seeing
    • L1 0382-0405
    • L2 0367-0391
  • UT 08:59 - next field GSPC 247U
    • seeing briefly improved to 1.2" but now back up to 1.5" - if we can stay at 1.2" will do StoneD L2 MOS
    • L1 0406-0429
    • L2 0392-0414

Rotator ZP Part Deux

  • Stone D is now up, will attempt that in between seeing bubbles
  • UT 09:25 - authorizing LUCI2 only
  • low elevation, i guess seeing in this part of sky is a bit wonkier.....
  • bad seeing, but can centroid (1.7" FWHM K-band)
    • x-shift =0.94", yshift = -3.27", theta = -0.071
    • err 0.096",0.136"
  • should probably redo, but consistent with Stone A mask
  • taking confirmatory image delayed, the Sci part of the script had StoneO not StaoneD in the mos mask - so waiting....
  • we already had on confirmatory image (no need for shifts) - so putting LUCI-2 in safe config and reconfiguring for LBCs
  • UT 09:52 - reconfiguring for LBCs

LBC Checkout

Photometric fields
  • UT 10:19 - co-pointing error again....
  • sdssY4_r015d14
  • UT 10:59 - seeing ~1.15" LBCR
  • UT 11:02 - moving to another photometric field
  • sdssY3_r2349d15
  • UT 11:31 - seeing ~1.5" on LBC images, DIMM ahowed 2.3" on last reading
2018 XG5
  • UT11:39 - sending collimation preset
  • UT 11:46 - DIMM shows 1.97"
  • UT 11:50 - starting science-like OB
  • UT 12:02 - LBC-R 115654 - wrap around problem - IT 8489 (affected image is lbcr.20220906.115540.fits too few/many downloaded pixels error -opk)
  • UT 12:06 - hit brightness limits LBCB 120146, LBCR 120027 near limit (56k counts)
  • UT 12:07 - stopping script - moving to skyflats - got 2 good LBCB, 2 good LBCR

LBC Twilight Flats

  • sloan g, sloan i PA=0
  • sloan g, sloan i PA=180
  • Uspec, F972 PA=0
  • Uspec, F972 PA=180

Closed Dome LUCI Imaging Flats (N375)

  • putting L2 mask away while telescope moves to park position
  • UT 12:50 - starting LUCI N375 imaging cals- still see saturn persistence (oops)
    • Y1 L1 0430-0439
    • Y1 L2 0420-0429
    • FeII L1 0440-0449
    • FeII L2 0430-0439
    • H2 L1 0450-0459
    • H2 L2 0440-0449
    • z L1 0460-0469
    • z L2 0450-0459
    • J L1 0470-0479
    • J L2 0460-0469
    • H L1 0480-0489
    • H L2 0470-0479
    • K L1 0490-0499
    • K L2 0480-0489
    • Ks L1 0500-0509
    • Ks L2 0490-0499
  • UT 13:35 - putting LUCIs in safe config

-- BarryRothberg - 05 Sep 2022
Topic revision: r7 - 07 Sep 2022, BarryRothberg
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback